OpenStudy (anonymous):

What is the product of 2 sqrt(10x) * 3sqrt(6x)

OpenStudy (lalaly):

\[2\sqrt{10x} \times 3\sqrt{6x} = 2 \times 3 (\sqrt{(6x \times 10x)}\]

OpenStudy (lalaly):

=\[6\sqrt{60x^2}= 6x \sqrt60\] and\[\sqrt 60 = \sqrt{4 \times 15} = 2\sqrt15\]so the whole answer would be\[6x \times 2 \sqrt15 = 12x \sqrt15\]

OpenStudy (anonymous):

ty lalay, outta all the others, your by far the best

OpenStudy (lalaly):

ur welcome, i cn help u wit ur other questions

OpenStudy (anonymous):

ok ty very much lana,